OpenStudy (anonymous):

add and simplify if possible 2/9z^3 + 1/6z^2

jimthompson5910 (jim_thompson5910):

\[\Large \frac{2}{9z^3} + \frac{1}{6z^2}\] \[\Large \frac{2*2}{2*9z^3} + \frac{1}{6z^2}\] \[\Large \frac{4}{18z^3} + \frac{1}{6z^2}\] \[\Large \frac{4}{18z^3} + \frac{3z*1}{3z*6z^2}\] \[\Large \frac{4}{18z^3} + \frac{3z}{18z^3}\] \[\Large \frac{4+3z}{18z^3}\] So \[\Large \frac{2}{9z^3} + \frac{1}{6z^2}=\frac{4+3z}{18z^3}\]